OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (anonymous):

Okay the only way to solve this is via the graphing method. Im too lazy to give you the full answer, but i will summarize for you. for Part 1, system with one solution, When you graph out a system of equations ( let say there is only 2). The intercept of the 2 lines, the point where they intercept is the solution. So if they intercept at only one point there is only one solution. For part 2, system with no solution, for 2 lines to not intersect, is when both lines are parallel. if both lines are parallel they will never intersect even when stretched to infinity. For part 3, system with infinite solution, it is when both lines intersect at every point, which means to say they are the same line An example will be x+y=2 2x+2y=4 Both are the same line when simplified.

OpenStudy (anonymous):

PLEASE DO FULL QUESTION LOL ALL I NEED IS EXAMPLES FOR THEM , FOR EACH ONE EXAMPLES

OpenStudy (anonymous):

Dont be lazy come on, i describe everything in detail for you. But oh well, since it is so easy... Part 1, y=x+2 y=2x+4.. Part 2, y=x+2 y=x+4 Part 3, y=x+1 2y=2x+2